A soap bubble (n=1.35) is floating in air . If the thickness of the bubble wall is 115 nm, dertemine the wavelength of the visible light that is most strongly reflected?


NOTE: NEEDED AS FAST AS POSSIBLE
1
Expert's answer
2020-07-01T15:41:34-0400

The condition is


"2nt=(m+1/2)\lambda."

Express the wavelength:


"\lambda=\frac{2nt}{m+1/2},\\\space\\
\lambda=\frac{2\cdot1.35\cdot115}{0+1/2}=621\text{ nm}."